OpenStudy (anonymous):

**WILL ALWAYS FAN+MEDAL** Using the Heron's Formula how do you find the area of a triangle(A,B,C) with 3 side lengths which are: A: 22 cm B: 17 cm C: 27 cm Can somebody tell me how to do this by any chance? ???

OpenStudy (anonymous):

@YanaSidlinskiy @YOUNGSUPERMAN @adajiamcneal @nincompoop @Anestasia @amistre64 @deana99 @abdullah13TEEN @AaronAndyson @Amelia_13 @january123 @adajiamcneal @Daniela_Leah @horsegirl27 @HufflepuffKlaine @THEHELPER123 @GreenPanda333 @em2000 @e.mccormick

OpenStudy (anonymous):

186.67619023324855 Square Units @HackberryAbby18

OpenStudy (anonymous):

Thanks by chance can you walk me through that? Sorry :P

OpenStudy (zarkon):

let \(s=(22+17+27)/2\) then compute \[\sqrt{s(s-22)(s-17)(s-27)}\]

OpenStudy (anonymous):

Thanks!! I wish I could medal both of you :P

OpenStudy (anonymous):

:) np

OpenStudy (zarkon):

I don't get 186.67619023324855

OpenStudy (anonymous):

??

OpenStudy (zarkon):

now i do...put in an extra 17 instead of 27

OpenStudy (zarkon):

\[132\sqrt{2}\]

OpenStudy (anonymous):

Lol Ok thanks lol

OpenStudy (anonymous):

@Zarkon can you fan me by chance you too @em2000

OpenStudy (anonymous):

the formula of a triangle is \[A=\frac{ 1 }{ 2 }\times base \times hieght\]

OpenStudy (e.mccormick):

@THEHELPER123 That does not help when you do not have the height, which is what the whole Heron's Formula part was about.